# Extracting the Speed of Sound in Heavy-Ion Collisions: A Study of Quantum Fluctuations and Thermalization

### v2

@@ Abstract
The thermalization of quark-gluon plasma created in heavy-ion collisions is crucial for understanding its behavior as a relativistic fluid and the thermodynamic properties of the Quantum Chromodynamics (QCD). This study investigates the role of fluctuations in the relationship between transverse momentum and particle multiplicity, with a particular focus on their impact on extracting the QCD speed of sound. In a thermalized quark-gluon plasma, fluctuations are dominantly quantum in origin and follow a Gaussian distribution due to their independence from the thermodynamic response. In contrast, non-thermalized systems display non-Gaussian fluctuations, reflecting the breakdown of thermalization. By leveraging the Gaussianity condition of quantum fluctuations, the physical value of the speed of sound can be extracted statistically, even in the presence of significant event-by-event fluctuations. This framework provides a robust diagnostic tool for probing thermalization and extracting thermodynamic properties in both large and small collision systems.
@@
Heavy-ion collisions at facilities such as the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) create extreme conditions, where Quark-Glasma Plasma (QGP) forms Shuryak (2017). A crucial question is whether this deconfined state of quarks and gluons achieves local thermal equilibrium, at least transiently, so that one is allowed to apply hydrodynamic models Gale et al. (2013); Romatschke and Romatschke (2019); Florkowski et al. (2018) and explore the phase structure of the strongly interacting matter through fundamental thermodynamic properties Luo and Xu (2017). As a closed quantum system of high energies, the thermalization of QGP is also essential for understanding non-equilibrium dynamics, in a broad sense Schlichting and Teaney (2019); Shen and Yan (2020); Berges et al. (2021); Brauner et al. (2022).
@@
Local thermalization of QGP in nucleus-nucleus collisions is widely accepted, supported by different types of flow signatures that reveal anisotropic particle emissions. These are indirect evidence, however, as they are consequences of hydrodynamic response of a fluid-like QGP with respect to the geometric deformation of initial state Heinz and Snellings (2013); Yan (2018). More explicitly, by assuming the QGP as a relativistic fluid, the thermodynamic quantities like pressure $P$, energy density $e$, and entropy $S$ can be well implemented in hydrodynamic models, and the initial geometric deformation results in anisotropies of pressure gradients, which further drives the system to expand anisotropically Ollitrault (1992). Although flow analyses have extended to the smaller colliding systems, including proton-lead Abelev et al. (2013); Aad et al. (2013); Khachatryan et al. (2015), proton-proton Khachatryan et al. (2010); Aad et al. (2016a); Acharya et al. (2024a) and even electron-positron Chen et al. (2024), aside from which, investigation of the local thermalization of the strongly interacting matter through directly the thermodynamic properties, such as the QCD speed of sound Van Hove (1982), was realized only recently Gardim et al. (2020a, b).
@@
In high-energy ultracentral collisions of heavy nuclei, where geometric deformation is substantially suppressed, if the created QGP reaches local thermal equilibrium, the QCD speed of sound governs the response of variation in the observed mean transverse momentum of particles, $\langle p_{T}\rangle\equiv\sum p_{T}/N_{\rm c}$, to the variation in the charged particle multiplicity, $N_{\rm c}$. This is because, in ultracentral collisions, mean transverse momentum measures effectively temperature in the system Gardim et al. (2020a), while the charged particle multiplicity is linked to the entropy. Consequently, linear relationships exist between $\langle p_{T}\rangle$ and $T$, and between $N_{\rm c}$ and $S$. Corresponding to Eq. (1), the response is purely thermodynamic and deterministic, so one has Gardim et al. (2020a, b)
@@
where $\langle p_{T}\rangle_{0}$ and $N_{0}$ are the averaged values of the mean transverse momentum and the charged particle multiplicity over the ultracentral collision events, while $\Delta_{p}=\langle p_{T}\rangle-\langle p_{T}\rangle_{0}$ and $\Delta_{N}=N_{\rm c}-N_{0}$ are variations. Note that the mechanism of relating $\Delta_{p}$ and $\Delta_{N}$ is not unique. From purely multi-parton interactions or mini-jet production, for instance, a linear response relation between $\Delta_{p}$ and $\Delta_{N}$ presents analogously, but the coefficient has of course, nothing to do with the speed of sound. Nonetheless, it is crucial that Eq. (2) from a thermalized QGP is thermodynamic, while the correlations induced in non-thermal systems are quantum.
Figure 1: The normalized two-dimensional joint probability distribution of $\mathcal{P}(\Delta_{N},\Delta_{p})$, for the central Pb-Pb (0-5%) and p-Pb (0-2%) collisions at $\sqrt{s_{NN}}=5.02$ TeV, from simulations using hydrodynamics, HIJING and PYTHIA. The growth of width of the distributions, from (a) to (d), indicating the increasing of fluctuation strength. The tilted tip of the distributions (except PYTHIA) implies a positive correlations between $\Delta_{p}$ and $\Delta_{N}$.
@@
On the other hand, there are fluctuations in realistic heavy-ion collisions. On an event-by-event basis, these fluctuations are dominantly of quantum origin, such as the quantum fluctuations of the nucleons populating inside nuclei as they collide, and decay and scatterings of hadrons during final-state particlization the . In terms of flow measurement, fluctuations in high-energy heavy-ion collisions have been well captured through multi-particle correlations Shen and Yan (2020), in consistent to hydrodynamic model simulations on an event-by-event basis. The role of fluctuations in the response relation between $\Delta_{p}$ and $\Delta_{N}$, in the ultracentral collisions, remains unclear so far Sun and Yan (2024). In this Letter, we will investigate the effect of fluctuations on the extraction of speed of sound, and we propose a statistical method that systematically isolates the thermodynamic response, so that in both theoretical simulations and experiments the speed of sound can be well identified in the presence of the fluctuations. More importantly, we find that the quantum nature of the event-by-event fluctuations makes the extraction of the speed of sound an ideal probe of the QGP thermalization.
@@
Quantum fluctuations. – In realistic heavy-ion collisions, with respect to the mean transverse momentum and the charged particle multiplicity, the presence of fluctuations leads to a two-dimensional joint probability distribution of $\mathcal{P}(\Delta_{p},\Delta_{N})$ Samanta et al. (2024, 2023); Aad et al. (2024), instead of the deterministic relation Eq. (2).
@@
While such a distribution can be measured experimentally, for the purpose of illustration, we generate a set of $\mathcal{P}(\Delta_{p},\Delta_{N})$ from numerical simulations of hydrodynamic model and non-thermal models num . The hydrodynamic model is hybrid that incorporates the generation of initial density profile using the TRENTo method Moreland et al. (2015), solving coupled equations of viscous hydrodynamics with the EOS from lattice QCD (LEOS) Bazavov et al. (2014) for the system’s dynamical evolution using the MUSIC program Schenke et al. (2011, 2010); Paquet et al. (2016), which is then connected to UrQMD for particlization and hadron gas evolution Bass et al. (1998). These are standard theoretical prescriptions that systematically takes into account the event-by-event quantumn fluctuations from different stages. It should be emphasized that while solving hydrodynamics implies local thermalization up to the dissipative corrections, the hadron dynamics regarding UrQMD does not require local thermalization. Therefore, local thermalization from the numerical simulations of the hybrid hydrodynamic models is only transiently realized.
@@
For lead-lead (Pb-Pb) collisions at high energies, longitudinal boost invariance is a good approximation, which allows us to simplify the analysis and work in a 2+1 dimensional setup. However, this longitudinal symmetry is broken in small collision systems such as proton-lead (p-Pb), for which we invoke the exact 3+1 dimensional characterization of the initial state in TRENTo-3D Soeder et al. (2023), as well as in MUSIC. For comparison, we also simulate and collect data on the mean transverse momentum and multiplicity in p-Pb collisions using non-thermal models, including HIJING Wang and Gyulassy (1991); Deng et al. (2011); Wang and Gyulassy (1992) and PYTHIA Sjostrand et al. (2008); Sjöstrand et al. (2015); Bierlich et al. (2022). In these non-thermal models, particles are generated by parton scatterings and radiative processes rather than thermal medium expansion, meaning that local thermalization is absent.
@@
Figure 2: The probability distribution $\rho(\delta|c_{s}^{2})$ with $c_{s}^{2}$ determined by the zero skewness condition $\{\delta^{3}\}_{c}=0$ (red histogram) and $\{\delta^{5}\}_{c}=0$ (blue histogram). For comparison, the Gaussian distributions with zero mean and the variance $\{\delta^{2}\}^{1/2}$ determined with respect to the extracted $c_{s}^{2}$ values are shown by lines of the same colors.
@@
With the $\langle p_{T}\rangle_{0}$ and $N_{0}$ fixed with respect to the top central collisions in experiments, for instance, top 5% centrality for Pb-Pb collisions at the LHC with respect to the pseudo-rapidity $|\eta|<0.5$ and particles of transverse momentum $p_{T}\in[0,3]$ GeV Aad et al. (2016b); Acharya et al. (2023), from numerical simulations we obtain the two-dimensional joint probability distributions, as presented in Fig. 1. In Pb-Pb collisions, the vanishing impact parameter, or saturation of system size, appears only in the ultracentral events, namely, the right tip of the distribution in Fig. 1 (a) Samanta et al. (2023); Gardim et al. (2020b). In p-Pb collisions, on the other hand, the geometry of the generated system is dominantly determined by the proton, one does not require the condition of ultracentral collisions to suppress the size dependence Gardim et al. (2024). In principle, the analysis of the speed of sound can be carried out through most centralities of the p-Pb collisions. Nonetheless, we still focus on the top 2% central events in p-Pb, where system collective expansion has been observed in experiments through particle flow of multi-particle correlations Khachatryan et al. (2015).
For later convenience, we have also obtained the effective temperatures from the freeze-out hypersurface Gardim et al. (2020a), with details given in the Supplemental Materials. While for the Pb-Pb system, one approximately has $\langle p_{T}\rangle\approx 3T_{\rm eff}$ due to the boost invariance, for p-Pb collisions, on an event-by-event basis we find that $\langle p_{T}\rangle\approx 2.22\sim 2.8\,T_{\rm eff}$ and $\langle p_{T}\rangle\approx 2.16\sim 2.58\,T_{\rm eff}$ for the collision energies $\sqrt{s_{NN}}=5.02$ TeV and $8.16$ TeV, respectively.
@@
Importantly, for a thermalized QGP, $\delta$ is independent of the thermodynamic response relation. This independence reflects the quantum origin of $\delta$, which introduces stochastic fluctuations uncorrelated with the deterministic macroscopic relationship. While $\delta$ is independent of the thermodynamic response, it may exhibit statistical correlations with individual observables like $\Delta_{N}$ or $\Delta_{p}$, arising from shared quantum or initial-state fluctuations. In a thermalized system, where $\delta$ varies randomly and independently across events, its distribution approaches Gaussianity by the Central Limit Theorem (CLT) Samanta et al. (2024, 2023). On the other hand, in a system without thermalization, the response relation itself (where the response coefficient is no longer the speed of sound) is quantum, hence $\delta$ is not independent from event to event and the $\delta$-distribution is generically non-Gaussian. This non-Gaussinity in $\delta$ in non-thermal systems can be conceptually understood as a consenquence of interactions among these fluctuations, through the quantum response relation. This is analogous to the non-Gaussianity observed in primordial quantum fluctuations, which signals interactions in the early universe Maldacena (2003).
@@
Irrespective of the condition of local thermalization, finite-size effects, proportional to $1/N_{0}$, introduce residual non-Gaussianity to the $\delta$-distribution. Specifically, corrections to higher-order moments of the $\delta$-distribution scale as powers of $1/N_{0}$, with the skewness scaling as $1/N_{0}^{3/2}$, and kurtosis scaling as $1/N_{0}^{2}$, etc. Thus, the previous discussions of Gaussianity in thermlized QGP and non-Gaussianity in non-thermal systems apply up to corrections of order $1/N_{0}^{a}$, where $a$ depends on the specific observable or moment. These finite-size effects are particularly relevant when $N_{0}$ is relatively small.
@@
To verify the analysis, we need to solve the distribution of $\delta$ with respect to the physical value of the speed of sound. Regarding Eq. (3), it is a mathematical problem: How to solve the constant parameter $c_{s}^{2}$, given two known distributions of $\Delta_{p}$ and $\Delta_{N}$, and one unknown distribution of $\delta$? In fact, Eq. (3) implicitly defines the conditional distribution, $\rho(\delta|c_{s}^{2})$. In the small $c_{s}^{2}$ limit, $\rho(\delta|c_{s}^{2})$ approaches the distribution of $\Delta_{p}/\langle p_{T}\rangle_{0}$, while for $c_{s}^{2}\to\infty$, the distribution of $\delta$ is equivalent to that of $-\Delta_{N}/N_{0}$.
[…]
@@
with the physical value of the speed of sound its real root. The coefficients of the equation correspond respectively, to the standardized skewness or mixed skewness of the joint probability distribution $\mathcal{P}(\Delta_{p},\Delta_{N})$, which are all measurable in realistic experiments. With the speed of sound extracted from the zero skewness condition, one is allowed to solve the exact probability distribution $\rho(\delta|c_{s}^{2})$. As shown in Fig. 2, to a good approximation within corrections of $1/N_{0}^{a}$, we find that $\delta$ from hydrodynamic simulations follow the Gaussian distribution, while from non-thermal simulations, fluctuations of $\delta$ exhibit apparently non-Gaussianity with heavy tails. We have also solved similarly, with respect to the condition that the fifth order cumulant $\{\delta^{5}\}_{c}=\{\delta^{5}\}-10\{\delta^{2}\}\{\delta^{3}\}=0$, which gives consistent results. Besides, for the values of $c_{s}^{2}$ other than the desired ones, we notice that even from hydrodynamic models $\rho(\delta|c_{s}^{2})$ becomes non-Gaussian.
@@
where {}$\{\ldots\}_{I}$ denotes event average in the sub-bin $I$. If $\delta$ and $\Delta_{N}$ are independent in the sub-bins, the speed of sound can be simply read off as the slope from the resulted line between $\{\Delta_{p}\}_{I}/\langle p_{T}\rangle_{0}$ and $\{\Delta_{N}\}_{I}/N_{0}$, otherwise, the slope gets correction from $\{\delta\}_{I}\neq 0$. In fact, the condition $\{\delta\}=0$ implies already in the sub-bins, $\{\delta\}_{I}=\alpha\{\Delta_{N}\}_{I}/N_{0}$, with $\alpha$ a constant. One may understand this linear relation as the leading order expansion of $\{\delta\}_{I}$ in terms of $\{\Delta_{N}\}_{I}/N_{0}$.
@@
$c_{s}^{2}$ sub-bin slope $\{\delta^{3}\}_{c}=0$$\{\delta^{5}\}_{c}=0$ LEOS
PbPb (Hydro, 5.02TeV) $0.123\pm 0.035$ $0.217\pm 0.032$ $0.216\pm 0.041$ 0.222-0.242
pPb (Hydro, 8.16 TeV) $0.176\pm 0.004$ $0.292\pm 0.013$ $0.287\pm 0.012$ 0.282-0.309
pPb (Hydro, 5.02 TeV) $0.197\pm 0.004$ $0.318\pm 0.011$ $0.313\pm 0.008$ 0.269-0.304
pPb (PYTHIA, 5.02 TeV) $-0.032\pm 0.002$ $1.178\pm 0.006$ $1.352\pm 0.019$ 0.227-0.278
pPb (HIJING, 5.02 TeV) $0.079\pm 0.003$ $1.104\pm 0.019$ $1.171\pm 0.053$ 0.206-0.271
Table 1: A summary of the extracted speed of sound from different methods for various model simulations.
@@
In Fig. 3 (a), using the sub-bin method, the linear correlations between $\{\langle p_{T}\rangle\}_{I}/\langle p_{T}\rangle_{0}$ and $\{N_{\rm c}\}_{I}/N_{0}$ from the different models and systems are shown. Except the results from PYTHIA, where the linear relation has a negative slope, all the model simulations results indicate a positive linear correlation between $\Delta_{p}$ and $\Delta_{N}$. Notably, from the HIJING model, although there is no local thermalization, one still finds a slope that is sizable in magnitude. However, even from hydrodynamic simulations, we find that the values of the slope are substantially small comparing to the physical values of $c_{s}^{2}$ from the LEOS expectations, suggesting a finite and negative contribution from $\{\delta\}_{I}$, which is shown in Fig. 3 (b).
@@
Corresponding to the slope of the sub-bin linearity, and the Gaussianity condition through zero skewness and $\{\delta^{5}\}_{c}=0$, we solve the physical values of the speed of sound . A summary of the obtained results is given in Table. 1. We notice that the extracted values of the speed of sound for all the hydrodynamic model simulations from the Gaussian distribution condition are compatible to the LEOS expectations, within uncertainties. For the non-thermal model simulations, however, the same procedure gives rise to non-physical values of the speed of sound, as they are greater than unity (speed of light $c=1$ as we take the natural units), which violates the causality condition Cherman et al. (2009); Hippert et al. (2025).

### v3

# Extracting the Speed of Sound in Heavy-Ion Collisions: A Study of Quantum-Initiated Fluctuations and Thermalization

@@ Abstract
The thermalization of quark-gluon plasma created in heavy-ion collisions is crucial for understanding its behavior as a relativistic fluid and the thermodynamic properties of the Quantum Chromodynamics (QCD). This study investigates the role of fluctuations in the relationship between transverse momentum and particle multiplicity, with a particular focus on their impact on extracting the QCD speed of sound. In a thermalized quark-gluon plasma, these fluctuations mostly originate from quantum fluctuations in the colliding nuclei, and exhibit a Gaussian distribution as a consequence of their independence from thermodynamic response. In contrast, non-thermalized systems display non-Gaussian fluctuations, reflecting the breakdown of thermalization. By leveraging the Gaussianity condition of quantum-initiated fluctuations, the physical value of the speed of sound can be extracted statistically, even in the presence of significant event-by-event fluctuations. This framework provides a robust diagnostic tool for probing thermalization and extracting thermodynamic properties in both large and small collision systems.
@@
Heavy-ion collisions at facilities such as the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) create extreme conditions, where Quark-Gluon Plasma (QGP) forms Shuryak (2017). A crucial question is whether this deconfined state of quarks and gluons achieves local thermal equilibrium, at least transiently, so that one is allowed to apply hydrodynamic models Gale et al. (2013); Romatschke and Romatschke (2019); Florkowski et al. (2018) and explore the phase structure of the strongly interacting matter through fundamental thermodynamic properties Luo and Xu (2017). As a closed quantum system of high energies, the thermalization of QGP is also essential for understanding non-equilibrium dynamics, in a broad sense Schlichting and Teaney (2019); Shen and Yan (2020); Berges et al. (2021); Brauner et al. (2022).
@@
Local thermalization of QGP in nucleus-nucleus collisions is widely assumed, based on different types of flow signatures that reveal anisotropic particle emissions. These are indirect evidence, however, as they are consequences of hydrodynamic response of a fluid-like QGP with respect to the geometric deformation of initial state Heinz and Snellings (2013); Yan (2018). More explicitly, by assuming the QGP as a relativistic fluid in which local thermalization is implied, the thermodynamic quantities like pressure $P$, energy density $e$, and entropy $S$ can be well implemented in hydrodynamic models, and the initial geometric deformation results in anisotropies of pressure gradients, which further drive the system to expand anisotropically Ollitrault (1992). The assumption of QGP local thermalization has been successfully applied to flow analyses in central and non-central nucleus-nucleus collisions, and has been extended to the smaller colliding systems, including proton-lead Abelev et al. (2013); Aad et al. (2013); Khachatryan et al. (2015), proton-proton Khachatryan et al. (2010); Aad et al. (2016a); Acharya et al. (2024a) and even electron-positron Chen et al. (2024). Nonetheless, investigation of local thermalization of QGP directly through the thermodynamic properties, such as the QCD speed of sound Van Hove (1982), was realized only recently Gardim et al. (2020a, b).
@@
In high-energy ultracentral collisions of heavy nuclei, where geometric deformation is substantially suppressed, if the created QGP reaches local thermal equilibrium, the QCD speed of sound governs the response of variation in the observed mean transverse momentum of particles, $\langle p_{T}\rangle\equiv\sum p_{T}/N_{\rm c}$, to the variation in the charged particle multiplicity, $N_{\rm c}$. This is because, in ultracentral collisions, mean transverse momentum measures the temperature in the system effectively Gardim et al. (2020a), while the charged particle multiplicity is linked to the entropy. Consequently, linear relationships exist between $\langle p_{T}\rangle$ and $T$, and between $N_{\rm c}$ and $S$. Corresponding to Eq. (1), the response is purely thermodynamic and deterministic, so one has Gardim et al. (2020a, b)
@@
where $\langle p_{T}\rangle_{0}$ and $N_{0}$ are the averaged values of the mean transverse momentum and the charged particle multiplicity over the ultracentral collision events, while $\Delta_{p}=\langle p_{T}\rangle-\langle p_{T}\rangle_{0}$ and $\Delta_{N}=N_{\rm c}-N_{0}$ are variations. Note that the mechanism of relating $\Delta_{p}$ and $\Delta_{N}$ is not unique. From purely multi-parton interactions or mini-jet production, for instance, a linear response relation between $\Delta_{p}$ and $\Delta_{N}$ presents analogously, but the coefficient has of course, nothing to do with the speed of sound.
Figure 1: The normalized two-dimensional joint probability distribution of $\mathcal{P}(\Delta_{N},\Delta_{p})$, for the central Pb-Pb (0-5%) and p-Pb (0-2%) collisions at $\sqrt{s_{NN}}=5.02$ TeV, from simulations using hydrodynamics, HIJING and PYTHIA. The tilted tip of the distributions (except PYTHIA) implies a positive correlation between $\Delta_{p}$ and $\Delta_{N}$.
@@
On the other hand, there are fluctuations in realistic heavy-ion collisions. On an event-by-event basis, these fluctuations are dominantly of quantum origin, such as the fluctuations of nucleons populating inside nuclei as they collide, and decay and scatterings of hadrons during final-state particlization the . In terms of flow measurement, fluctuations in high-energy heavy-ion collisions have been well captured through multi-particle correlations Shen and Yan (2020), consistent with hydrodynamic model simulations on an event-by-event basis. The role of fluctuations in the response relation between $\Delta_{p}$ and $\Delta_{N}$, in the ultracentral collisions, remains unclear so far Sun and Yan (2025). In this Letter, we will investigate the effect of fluctuations on the extraction of speed of sound, and we propose a statistical method that systematically isolates the thermodynamic response, so that in both theoretical simulations and experiments the speed of sound can be well identified in the presence of the fluctuations. More importantly, we find that these quantum-initiated event-by-event fluctuations qua make the extraction of the speed of sound an ideal probe of the QGP thermalization.
@@
Quantum-initiated fluctuations. – In realistic heavy-ion collisions, with respect to the mean transverse momentum and the charged particle multiplicity, the presence of fluctuations leads to a two-dimensional joint probability distribution of $\mathcal{P}(\Delta_{p},\Delta_{N})$ Giacalone et al. (2021); Samanta et al. (2024, 2023); Aad et al. (2024), instead of the deterministic relation Eq. (2).
@@
While such a distribution can be measured experimentally, for the purpose of illustration, we generate a set of $\mathcal{P}(\Delta_{p},\Delta_{N})$ from numerical simulations of hydrodynamic model and non-thermal models num . The hydrodynamic model is hybrid that incorporates the generation of initial density profile using the TRENTo method Moreland et al. (2015), solving coupled equations of viscous hydrodynamics with the EOS from lattice QCD (LEOS) Bazavov et al. (2014) for the system’s dynamical evolution using the MUSIC program Schenke et al. (2011, 2010); Paquet et al. (2016); hyd , which is then connected to UrQMD for particlization and hadron gas evolution Bass et al. (1998). By construction, the quantum-initiated fluctuations that are dominantly from the initial state are accounted for stochastically in the hybrid modeling. It should be emphasized that while solving hydrodynamics implies local thermalization up to the dissipative corrections, the hadron dynamics regarding UrQMD does not require local thermalization. Therefore, local thermalization in the hybrid hydrodynamic model is only transiently realized.
@@
For lead-lead (Pb-Pb) collisions at high energies, longitudinal boost invariance is a good approximation, which allows us to simplify the analysis and work in a 2+1 dimensional setup. However, this longitudinal symmetry is broken in asymmetric systems such as proton-lead (p-Pb), for which we invoke the exact 3+1 dimensional characterization of the initial state in TRENTo-3D Soeder et al. (2023), as well as in MUSIC. For comparison, we also simulate and collect data on the mean transverse momentum and multiplicity in p-Pb collisions using non-thermal models, including HIJING Wang and Gyulassy (1991); Deng et al. (2011); Wang and Gyulassy (1992) and PYTHIA Sjostrand et al. (2008); Sjöstrand et al. (2015); Bierlich et al. (2022). In these non-thermal models, particles are generated by parton scatterings and radiative processes rather than thermal medium expansion, without local thermalization.
@@
Figure 2: The normalized probability distribution $\rho(\delta|c_{s}^{2})$, with $c_{s}^{2}$ determined by the zero skewness condition $\{\delta^{3}\}_{c}=0$ (red solid histogram). For comparison, the Gaussian distributions with zero mean and the variance $\{\delta^{2}\}^{1/2}$ determined with respect to the extracted $c_{s}^{2}$ values are shown by red solid line.
@@
With the event averaged values of $\langle p_{T}\rangle$ and $N_{c}$ fixed with respect to the top 5% centralities for Pb-Pb collisions and top 2% centralities for p-Pb collisions at the LHC energies, for particles of pseudo-rapidity $|\eta|<0.5$ and transverse momentum $p_{T}\in[0,3]$ GeV Aad et al. (2016b); Acharya et al. (2023); Khachatryan et al. (2015), from numerical simulations we obtain the two-dimensional joint probability distributions, as presented in Fig. 1. Note that from the hydrodynamic simulations of central Pb-Pb collisions (panel (a)), to HIJING and PYTHIA simulations of central p-Pb collisions (panel (c) and (d)), the width of these distributions grows significantly, indicating increasing fluctuation strength. In Pb-Pb collisions, the vanishing impact parameter, or saturation of system size, appears only in the ultra-central events, namely, the right tip of the distribution in Fig. 1 (a) Samanta et al. (2023); Gardim et al. (2020b). In p-Pb collisions, although initial geometry is dominantly determined by the proton, the effective system size saturates only in the ultra-central collisions Gardim et al. (2024). Therefore, in the following for the extraction of speed of sound, we only focus on the 0-$2.5\%$ centrality class of Pb-Pb collisions and 0-1% centrality class in p-Pb collisions. For later convenience, we also obtained the effective temperatures from the freeze-out hypersurface Gardim et al. (2020a); Sup ,.
@@
Importantly, for a thermalized QGP, $\delta$ is independent of the thermodynamic response relation. This independence reflects the quantum origin of $\delta$, which introduces stochastic fluctuations uncorrelated with the deterministic relationship. While $\delta$ is independent of the thermodynamic response, it may exhibit statistical correlations with individual observables like $\Delta_{N}$ or $\Delta_{p}$, arising from shared quantum or initial-state fluctuations. In a thermalized system, where $\delta$ varies randomly and independently across events, its distribution approaches Gaussianity by the Central Limit Theorem (CLT) Samanta et al. (2024, 2023). On the other hand, in a system without thermalization, the response relation itself (where the response coefficient is no longer the speed of sound) is quantum, hence $\delta$ is not independent from event to event and the $\delta$-distribution is generically non-Gaussian. This non-Gaussianity in $\delta$ in non-thermal systems can be conceptually understood as a consequence of interactions among these fluctuations, through the quantum response relation. This is analogous to the non-Gaussianity observed in primordial quantum fluctuations, which signals interactions in the early universe Maldacena (2003).
@@
Irrespective of the condition of local thermalization, finite-size effects, proportional to $1/N_{0}$, introduce residual non-Gaussianity to the $\delta$-distribution. Specifically, corrections to higher-order moments of the $\delta$-distribution scale as powers of $1/N_{0}$, with the skewness scaling as $1/N_{0}^{3/2}$, and kurtosis scaling as $1/N_{0}^{2}$, etc. Thus, the previous discussions of Gaussianity in thermalized QGP and non-Gaussianity in non-thermal systems apply up to corrections of order $1/N_{0}^{a}$, where $a$ depends on the specific observable or moment. These finite-size effects are particularly relevant when $N_{0}$ is relatively small.
@@
To verify the analysis, we need to solve the distribution of $\delta$ with respect to the physical value of the speed of sound. Regarding Eq. (3), it is a mathematical problem: How to solve the constant parameter $c_{s}^{2}$, given two known distributions of $\Delta_{p}$ and $\Delta_{N}$, and one unknown distribution of $\delta$? In fact, Eq. (3) implicitly defines the conditional distribution, $\rho(\delta|c_{s}^{2})$. In the small $c_{s}^{2}$ limit, $\rho(\delta|c_{s}^{2})$ approaches the distribution of $\Delta_{p}/\langle p_{T}\rangle_{0}$, while for $c_{s}^{2}\to\infty$, the distribution of $\delta$ is equivalent to that of $\Delta_{N}/N_{0}$.
One first notices a trivial relation, by definition, in the ultracentral collisions, $\{\delta\}=\frac{1}{N_{\rm events}}\sum_{\rm events}\delta=0$, where the brackets $\{\ldots\}$ denote average over the ultracentral events. For a thermalized QGP, given the joint probability distribution $\mathcal{P}(\Delta_{p},\Delta_{N})$ with the physical value of the speed of sound, the fluctuation $\delta$ is Gaussian. Consequently, the physical value of the speed of sound implies vanishing cumulants of the $\delta$-distribution of arbitrary orders, up to corrections of $1/N_{0}^{a}$. To proceed, the simplest calculation is to solve the zero skewness condition, $\{\delta^{3}\}_{c}=\{\delta^{3}\}=0$, and then we examine the Gaussianity of the $\delta$-distribution accordingly. The zero skewness condition yields an equation of $c_{s}^{2}$, namely,
@@
with the physical value of the speed of sound its real root. The coefficients of the equation correspond respectively, to the standardized skewness or mixed skewness of the joint probability distribution $\mathcal{P}(\Delta_{p},\Delta_{N})$, which are all measurable in realistic experiments. With the speed of sound extracted from the zero skewness condition, one is allowed to solve the exact probability distribution $\rho(\delta|c_{s}^{2})$. As shown in Fig. 2, to a good approximation within corrections of $1/N_{0}^{a}$, we find that $\delta$ from hydrodynamic simulations follow the Gaussian distribution, while from non-thermal simulations, fluctuations of $\delta$ exhibit apparently non-Gaussianity with heavy tails. We have also solved similarly, with respect to the condition that the fifth order cumulant $\{\delta^{5}\}_{c}=\{\delta^{5}\}-10\{\delta^{2}\}\{\delta^{3}\}=0$, which gives consistent results Sup . Besides, for the values of $c_{s}^{2}$ other than the desired ones, we notice that even from hydrodynamic models $\rho(\delta|c_{s}^{2})$ becomes non-Gaussian.
@@
where $\{\ldots\}_{I}$ denotes event average in the sub-bin II. If $\delta$ and $\Delta_{N}$ are independent in the sub-bins, the speed of sound can be simply read off as the slope from the resulted line between $\{\Delta_{p}\}_{I}/\langle p_{T}\rangle_{0}$ and $\{\Delta_{N}\}_{I}/N_{0}$, otherwise, the slope gets correction from $\{\delta\}_{I}\neq 0$. In fact, the condition $\{\delta\}=0$ implies already in the sub-bins, $\{\delta\}_{I}=\alpha\{\Delta_{N}\}_{I}/N_{0}$, with $\alpha$ a constant. One may understand this linear relation as the leading order expansion of $\{\delta\}_{I}$ in terms of $\{\Delta_{N}\}_{I}/N_{0}$.
@@
$c_{s}^{2}$ sub-bin slope $\{\delta^{3}\}_{c}=0$ $\{\delta^{5}\}_{c}=0$ LEOS $T_{\rm eff}$ (MeV)
PbPb (Hydro, 5.02TeV) $0.123\pm 0.035$ $0.217\pm 0.032$ $0.216\pm 0.041$ 0.222-0.242 210.3
pPb (Hydro, 8.16 TeV) $0.176\pm 0.004$ $0.292\pm 0.013$ $0.287\pm 0.012$ 0.282-0.309 319.9-382.1
pPb (Hydro, 5.02 TeV) $0.197\pm 0.004$ $0.318\pm 0.011$ $0.313\pm 0.008$ 0.269-0.304 280.2-353.4
pPb (PYTHIA, 5.02 TeV) $0.032\pm 0.002$ $1.178\pm 0.006$ $1.352\pm 0.019$ 0.227-0.278 203.2-256.4
pPb (HIJING, 5.02 TeV) $0.079\pm 0.003$ $104\pm 0.019 1.1$$171\pm 0.053$ 0.206-0.271 219.8-277.2
Table 1: A summary of the extracted speed of sound from different methods for various model simulations. By determining the effective temperatures via their linear relation with $\langle p_{T}\rangle_{0}$, $c_{s}^{2}$ from LEOS can be estimated Bazavov et al. (2014). Note that unlike in PbPb collisions where one approximately has $\langle p_{T}\rangle\approx 3T_{\rm eff}$, in pPb collisions longitudinal expansion introduces larger uncertainty in the proportionality coefficient.
@@
In Fig. 3 (a), using the sub-bin method, the linear correlations between $\{\langle p_{T}\rangle\}_{I}/\langle p_{T}\rangle_{0}$ and $\{N_{\rm c}\}_{I}/N_{0}$ from the different models and systems are shown. Except the results from PYTHIA, where the linear relation has a negative slope, all the model simulations indicate a positive linear correlation between $\Delta_{p}$ and $\Delta_{N}$. Notably, from the HIJING model, although there is no local thermalization, one still finds a slope that is sizable in magnitude. However, even from hydrodynamic simulations, we find that the values of the slope are substantially small comparing to the physical values of $c_{s}^{2}$ from the LEOS expectations, suggesting a finite and negative contribution from $\{\delta\}_{I}$. Indeed, as demonstrated in Fig. 3 (b), where the averaged fluctuations in each sub-bin is plotted versus $N_{c}/N_{0}$, a clear negative correlation emerges between $\delta$ and $N_{c}$, reflecting the suppression of quantum fluctuations by larger system entropy.
@@
Corresponding to the slope of the sub-bin linearity, and the statistical method using Gaussianity condition, we solve the physical values of the speed of sound. A summary of the obtained results is given in Table. 1. We notice that the extracted values of the speed of sound for all the hydrodynamic model simulations from the Gaussian distribution condition are compatible to the LEOS expectations, within uncertainties. For the non-thermal model simulations, however, the same procedure yields non-physical values, as the speed of sound appears greater than unity (speed of light $c=1$ as we take the natural units), violating the causality condition Cherman et al. (2009); Hippert et al. (2025). Furthermore, since the extraction of $c_{s}^{2}$ via the linear slopes of $\Delta_{p}$ and $\Delta_{N}$ is systematically biased by the negative correlation between $\delta$ and $N_{c}$, determining $c_{s}^{2}$ within the Gaussianity framework is more reliable where such correlations are inherently accounted for.